# BSTcontrol Bimorph Lithium Niobate Piezoelectric Micromachined Ultrasonic Transducers

### v1

@@ Abstract
Piezoelectric micromachined ultrasonic transducers (PMUTs) are widely used in applications that demand mechanical resilience, thermal stability, and compact form factors. Lead zirconate titanate (PZT) and aluminum nitride (AlN) active layers are used in PMUTs to enable acoustic actuation, sensing, or bidirectional operation. These platforms rely on bimorph films to maximize electromechanical coupling () through thin-film deposition, which uses intermediate electrode layers to establish opposing electric fields. Consequently, incumbent PMUT platforms are limited in achievable film thickness and feature material interfaces that compromise mechanical integrity and thermal performance. Combined with the intrinsic limitations of PZT and AlN, these factors motivate exploration of alternative PMUT material platforms. Recent efforts have sought to demonstrate that single-crystal lithium niobate (LN) is a promising candidate, offering substantially higher and bidirectional performance. Advances in LN film transfer technology have enabled the formation of periodically poled piezoelectric (P3F) LN, facilitating a bimorph stack without intermediate electrodes. In this work, we showcase bimorph PMUTs incorporating a mechanically robust, 20 m thick P3F LN active layer. We establish the motivation for LN PMUTs through a material comparison, followed by extensive membrane geometry optimization and subsequent enhancement of the PMUT’s . We demonstrate a 775 kHz flexural mode device with a quality factor () of 200 and an extracted of 6.4%, yielding a high transmit efficiency of 65 nm/V with a mechanically robust active layer. We leverage the high performance to demonstrate extreme-temperature resilience, showcasing stable device operation up to 600 ∘C and survival up to 900 ∘C, highlighting LN’s potential as a resilient PMUT platform.
@@ 1 Introduction
PZT, a lead-based perovskite, remains dominant in the industry owing to its strong transmit characteristics [13]. However, its high electrical permittivity at constant stress () and the high associated dielectric loss limit sensing performance, such as the signal-to-noise ratio [14]. Despite these drawbacks, the maturity of PZT fabrication processes continues to sustain its role as a primary PMUT material. To overcome limitations faced by PZT, AlN is employed for its low dielectric loss and high sensitivity. Although it possesses moderate piezoelectricity, leading to moderate , its low and low dielectric loss enhance sensitivity and signal-to-noise ratio, making it ideal for low-power, high-sensitivity systems [15, 16]. Nevertheless, PZT and AlN exhibit complementary limitations: PZT excels as an actuator but performs poorly as a sensor, whereas AlN offers high sensitivity but limited actuation capability. A comparison is listed in Table 1 [17, 18, 19, 20, 21].
@@
| Material | Excita-tion | (pm/V) | (V/m/Pa) | ||
| PZT-5A | TFE | -171 | -0.011 | 1700 | 350 ∘C |
| KNN | TFE | -140 | -0.008 | 300 | 235 ∘C |
| ScAlN | TFE | 8.9 | 0.030 | 24.4 | 1100 ∘C |
| 36∘Y LN | TFE | -16.9 | -0.024 | 38.6 | 1200 ∘C |
| 36∘Y LN | LFE | 38.5 | 0.053 | 34.4 | 1200 ∘C |
| 128∘Y LN | TFE | -27.0 | -0.033 | 34.6 | 1200 ∘C |
| 128∘Y LN | LFE | -37.4 | -0.053 | 38.2 | 1200 ∘C |
| Y-cut LN | TFE | -20.3 | -0.027 | 43.6 | 1200 ∘C |
| Y-cut LN | LFE | 27.1 | 0.040 | 29.2 | 1200 ∘C |
| X-cut LN | LFE | -40.0 | -0.054 | 43.6 | 1200 ∘C |
@@
Single-crystal piezoelectric lithium niobate (LN) has recently emerged as a promising candidate to address these limitations [26, 27]. LN offers a balanced combination of high piezoelectric coupling coefficients and a moderate , situating it between KNN and ScAlN [28, 29, 30]. Lu et al. first demonstrated a unimorph lateral-field-excited (LFE) acoustic transducer with a silicon oxide (SiO2) passive layer, achieving a high of 4.5% and illustrating LN’s potential for efficient acoustic transduction [26]. Importantly, unlike KNN and ScAlN, LN is primarily available as a bulk single-crystal, allowing for precise thickness control via smart cut or controlled chemical-mechanical polishing (CMP) [31, 32]. Moreover, recent advances in LN-LN bonding technology now enable the production of high-quality bimorphs using periodically poled piezoelectric layers (P3F) [Fig. 1 (c)], which do not necessitate inter-layer electrodes, thereby expanding design flexibility and fabrication reliability [33, 34, 35]. These developments motivate a revisit of the LN PMUT.
P3F LN differs from conventional piezoelectric platforms in its ability to efficiently excite the flexural mode without the need for additional electrodes between active layers. Due to its anisotropy, the LN wafer orientation can be flipped during the LN-LN wafer bonding process to change the polarity of the appropriate piezoelectric coupling constant, effectively forming a bimorph stack without the need for electrical fields of opposite polarity. A theoretical framework regarding multi-layer P3F structures in anisotropic materials can be found in [36]. While P3F LN has been predominantly used in radio frequency (RF) acoustic applications, efforts to leverage its benefits for ultrasonic transduction have achieved significant success [27, 37, 38]. We aim to further develop this platform by providing a comprehensive theoretical and experimental framework around LN PMUT design and its capabilities.
@@ 2 Design and Simulation
To justify the selection of X-cut LN for acoustic transduction, material FoMs were compared across incumbent piezoelectric platforms (Table 1). The primary FoMs were the strain-charge piezoelectric coupling tensor coefficient (, in m/V), and the strain-voltage piezoelectric coupling tensor coefficient (, defined as in V/m/Pa), quantifying the mechanical strain generation due to an electrical input, and electrical field generation due to a mechanical input, respectively [40]. While is widely adopted to evaluate the transmission efficiency, characterizing receiving sensitivity depends on the measurement configuration. Often, the stress-charge piezoelectric coupling coefficient (, in C/m) is preferred over for materials with large capacitance densities, making charge read-out the optimum interfacing method with the device. However, LN’s large and low optimize the platform for voltage sensitivity. Consequently, we utilized and as transmit and receive FoMs, respectively. The open-circuit electrical permittivity and Curie temperature () served as secondary metrics to evaluate capacitance density and robustness.
The comparison made in Table 1 highlights key performance trade-offs between PZT-5A, KNN, 36% ScAlN, and LN, motivating our choice of piezoelectric material. PZT-5A demonstrates the highest relevant coupling coefficient, but possesses a large and a low , limiting sensing performance and resilience. KNN offers a similar transmission efficiency with a comparatively lower , but remains sub-optimal compared with ScAlN and LN. ScAlN achieves high and , yet it falls short of PZT and KNN in terms of transmit FoM. In contrast, different cuts of LN exhibit a very high , paired with a moderate , and a high , which motivates its use for robust, high-temperature PMUTs. Moreover, LN offers unprecedented design flexibility, supporting both thickness-field-excited (TFE) and LFE electrode configurations.
@@
The chosen diaphragm geometry introduces new key design parameters in addition to those reported by Lu et al., shown in Fig. 2 (b) [26]. Namely , and . is defined as the ratio of the ellipse’s minor axis (), to its major axis (), i.e., = . The parameter defines the fraction of the major axis covered by the electrodes, optimized to match the transverse energy distribution in the diaphragm while also providing space for routing to contact pads.
@@ 3.3 Measurements and Post-processing
To address performance detriments, the packaged device was thermally annealed in open air using a Silicon Nitride (SiN) resistive heater. The temperature was ramped at approximately 25 ∘C per minute to 400 ∘C, held for one hour, and then ramped back down to room temperature. The anneal was performed in an open-air environment due to potential contamination issues stemming from the organic epoxy used in the package.
@@
Following measurement validation and post-processing, the LN PMUT demonstrated high performance. Despite a lower resonance frequency and smaller than predicted, the device exhibited excellent efficiency. Given its mechanically robust geometry, the packaged device and the SiN heater were utilized to observe the device performance at elevated temperatures.
@@ 4.2 Receiving Performance
For the measured 65 nm/V peak actuation displacement at a of 200, a of 160 pm/Pa is simulated. Accordingly, the resulting open-circuit sensitivity () is 2.46 mV/Pa at 775 kHz. To validate the calculated value, we compare it to the open-circuit voltage obtained directly from the FEA, shown in Fig. 21. The simulation confirms the calculated estimate, showing 2.4 mV/Pa at 775 kHz. Additionally, Fig. 21 shows that the device achieves a peak open-circuit voltage at its parallel resonance, reaching up to 38 mV/Pa. As in the case of transmit efficiency, we decouple sensitivity from by utilizing the normalized open-circuit sensitivity (). Hence, we obtain a of 12 V/Pa. Nonetheless, sensitivity is highly influenced by the interfacing circuitry, and its effect is most pronounced at the system level, motivating further studies of bimorph LN-based sensors.
While external factors play a large role in receiving performance, a comparison with incumbent PMUT platforms highlights the LN PMUT’s high voltage sensitivity. State-of-the-art 36% ScAlN-based air-coupled PMUT shows an of 1 mV/Pa at 60 kHz via a pulse-echo measurement [48]. On the other hand, its air-coupled PZT counterpart for long-range detection shows 2 mV/Pa at 40 kHz through a pitch-catch experiment [1]. The performance of these ferroelectric platforms is further enhanced by incorporating a DC bias. For example, the sensitivity in a 15% ScAlN PMUT is increased from 6.43 mV/Pa in air at 150 kHz to 9.67 mV/Pa with an 80 V bias [49]. As an alternative, PMUTs can also utilize a dual-electrode structure to improve receiving performance. For instance, [50] shows 4.4 mV/Pa at 55 kHz in 30% ScAlN. By comparison, the bimorph LFE LN PMUT achieves a comparable while operating at a significantly higher frequency, showing potential for a highly sensitive, robust PMUT platform.
@@ 4.3 Remaining Limitations and Future Development
The remaining parasitic feedthrough observed in the post-anneal electrical measurements (Fig. 14) is critical for future LN sensor development. Although LN intrinsically exhibits low dielectric loss, parasitic feedthrough acts as a major bottleneck, limiting achievable performance. While this effect was partially mitigated with the anneal, residual parasitics constrain usable sensitivity, particularly near the parallel resonance. While thermal treatment appears beneficial, the fundamental origin of the feedthrough remains unresolved and warrants further studies.
@@ Acknowledgment
The authors would like to thank Dr. Todd Bauer and Dr. Kwok-Keung (KK) Law for helpful discussions. The authors would also like to thank Dr. Bichoy Bahr, Dr. Udit Rawat, and Dr. Yao Yu from the Texas Instruments Kilby Labs for valuable feedback.

### v2

# Bimorph Lithium Niobate Piezoelectric Micromachined Ultrasonic Transducers

@@ Abstract
Piezoelectric micromachined ultrasonic transducers (PMUTs) are widely utilized in applications that demand mechanical resilience, thermal stability, and compact form factors. Recent efforts have sought to demonstrate that single-crystal lithium niobate (LN) is a promising PMUT material platform, offering high electromechanical coupling () and bidirectional performance. In addition, advances in LN film transfer technology have enabled high quality periodically poled piezoelectric films (P3F), facilitating a bimorph piezoelectric stack without intermediate electrodes. In this work, we showcase a bimorph PMUT incorporating a mechanically robust, 20 m thick P3F LN active layer. We establish the motivation for LN PMUTs through a material comparison, followed by extensive membrane geometry optimization and subsequent enhancement of the PMUT’s . We demonstrate a 775 kHz flexural mode device with a quality factor () of 200 and an extracted of 6.4%, yielding a high transmit efficiency of 65 nm/V with a mechanically robust active layer. We leverage the high performance to demonstrate extreme-temperature resilience, showcasing stable device operation up to 600 ∘C and survival up to 900 ∘C, highlighting LN’s potential as a resilient PMUT platform.
@@ 1 Introduction
PZT, a lead-based perovskite, remains dominant in the industry owing to its strong transmit characteristics [13]. However, its high electrical permittivity at constant stress () and the high associated dielectric loss limit sensing performance, such as the signal-to-noise ratio [14]. Despite these drawbacks, the maturity of PZT fabrication processes continues to sustain its role as a primary PMUT material. To overcome the limitations of PZT, AlN is employed for its low dielectric loss and high sensitivity. Although it possesses moderate piezoelectricity, leading to moderate , its low and low dielectric loss enhance sensitivity and signal-to-noise ratio, making it ideal for low-power, high-sensitivity systems [15, 16]. Nevertheless, PZT and AlN exhibit complementary limitations: PZT excels as an actuator but performs poorly as a sensor, whereas AlN offers high sensitivity but limited actuation. A comparison is listed in Tables 1 and 2 [17, 18, 19, 20, 21].
@@
| Material | Excita-tion | (pm/V) | (V/m/Pa) | (C/m2) | ||
| PZT-5A | TFE | -171 | -0.011 | -5.4 | 1700 | 350 ∘C |
| KNN | TFE | -140 | -0.008 | -11.2 | 2000 | 235 ∘C |
| ScAlN | TFE | 8.9 | 0.030 | -0.7 | 33.64 | 1100 ∘C |
| 36∘Y LN | TFE | -16.9 | -0.024 | -1.8 | 65.2 | 1200 ∘C |
| 36∘Y LN | LFE | 38.5 | 0.054 | 4.6 | 67.2 | 1200 ∘C |
| 128∘Y LN | TFE | -27.0 | -0.033 | -2.7 | 50.5 | 1200 ∘C |
| 128∘Y LN | LFE | -37.4 | -0.053 | -4.5 | 63.5 | 1200 ∘C |
| Y-cut LN | TFE | -20.3 | -0.027 | -2.4 | 84.0 | 1200 ∘C |
| Y-cut LN | LFE | 27.1 | 0.040 | 3.2 | 58.3 | 1200 ∘C |
| X-cut LN | LFE | -40.0 | -0.054 | -4.6 | 69.7 | 1200 ∘C |
@@
| Material | Excitation | (C/m2) | / (GV/m) | |
| Epitaxial PZT | TFE | -24.0 | 308 | -8.8 |
| Sol-gel PZT | TFE | -17.7 | 1076 | -1.9 |
| Sputtered KNN | TFE | -10.0 | 260 | -4.3 |
| Sputtered ScAlN | TFE | -2.4 | 24.4 | -12.5 |
| Transferred X-cut LN | LFE | -5.3 | 41.1 | -14.6 |
*The subscripts are for TFE excitation and for LFE excitation.
Single-crystal piezoelectric lithium niobate (LN) has recently emerged as a promising candidate to address these limitations [26, 27]. LN offers a balanced combination of high piezoelectric coupling coefficients and a moderate , situating it between KNN and ScAlN [28, 29, 30, 31, 32]. Lu et al. first demonstrated a unimorph lateral-field-excited (LFE) acoustic transducer with a silicon oxide (SiO2) passive layer, achieving a high of 4.5% and illustrating LN’s potential for efficient acoustic transduction [26]. Importantly, unlike KNN and ScAlN, LN is primarily available as a bulk single-crystal, allowing for precise thickness control via smart cut or controlled chemical-mechanical polishing (CMP) [33, 34]. Moreover, recent advances in LN-LN bonding technology now enable the production of high-quality bimorphs using periodically poled piezoelectric layers (P3F) [Fig. 1 (c)], which do not necessitate inter-layer electrodes, thereby expanding design flexibility and fabrication reliability [35, 36, 37]. These developments motivate a revisit of the LN PMUT.
P3F LN differs from conventional piezoelectric platforms in its ability to efficiently excite the flexural mode without the need for additional electrodes between active layers. Due to its anisotropy, the LN wafer orientation can be flipped during the LN-LN wafer bonding process to change the polarity of the appropriate piezoelectric coupling constant, effectively forming a bimorph stack without the need for electrical fields of opposite polarity. A theoretical framework regarding multi-layer P3F structures in anisotropic materials can be found in [38]. While P3F LN has been predominantly used in radio frequency (RF) acoustic applications, efforts to leverage its benefits for ultrasonic transduction have achieved significant success [39, 40, 27, 41, 42]. We aim to further develop this platform by providing a comprehensive theoretical and experimental framework around LN PMUT design and its capabilities.
@@ 2 Design and Simulation
To justify the selection of X-cut LN for acoustic transduction, material FoMs were compared across incumbent piezoelectric platforms (Table 1). The primary metrics evaluated the material’s piezoelectric response under different electrical and mechanical boundary conditions. The strain-charge piezoelectric coupling tensor (, in m/V), and the strain-voltage piezoelectric coupling tensor (, defined as in V/m/Pa) were used to quantify mechanical strain generation due to an electrical input, and electrical field generation due to a mechanical input, respectively [44]. While these tensors describe the material when it is free to deform, the stress-charge piezoelectric coupling tensor (, in C/m2) relates the electrical field with stress when the material is clamped. Accordingly, and are widely adopted to evaluate material’s transmit efficiency, depending on physical boundary conditions. On the other hand, characterizing receiving sensitivity depends on the measurement configuration. Often, is preferred over for materials with large capacitance densities, making charge read-out the optimum interfacing method with the device. As a result, , , and were utilized as FoMs to characterize the material’s piezoelectric response. The relative () and Curie temperature () served as secondary metrics to evaluate capacitance density and robustness.
The comparison in Table 1 highlights key performance trade-offs among the bulk material constants of PZT-5A, KNN, 36% ScAlN, and LN, motivating our choice of piezoelectric material. PZT-5A demonstrates the highest coupling tensor coefficient, but possesses a large and a low , limiting sensing performance and resilience. KNN features a comparable and the highest coefficients, but remains sub-optimal compared with ScAlN and LN due to its high . ScAlN achieves high and , but it falls short of PZT and KNN in transmit FoMs. In contrast, different cuts of LN exhibit a very high , paired with a moderate , and a high , which motivates its use for robust, high-temperature PMUTs. Moreover, LN offers unprecedented design flexibility, supporting both thickness-field-excited (TFE) and LFE electrode configurations.
The bulk piezoelectric constants seen in Table 1 are significantly affected when a carrier substrate is introduced [45]. In this configuration, the material is clamped in-plane but free to deform out-of-plane, allowing additional tensor components to contribute to the measured piezoelectric response. Namely, an effective stress-charge piezoelectric coupling constant (, in C/m2) and an effective permittivity () can be defined. Their ratio (/, in GV/m) describes the generated voltage in the constrained film, acting as a receive FoM. Several material platforms, such as PZT, KNN, and ScAlN, exhibit substantial enhancement of their effective piezoelectric coefficients under in-plane clamping, as summarized in Table 2 [1, 46, 7, 47]. For example, the reported of -2.4 C/m2 in 36% ScAlN is markedly larger than its pure of only -0.7 C/m2, which was experimentally verified in [47]. To compare with these measurements, the effective material constants for LFE PMUTs in X-cut LN were derived from the stress-charge piezoelectric constitutive system while assuming zero in-plane strain and zero out-of-plane stress. The resulting parameters, shown in Table 2, indicate that X-cut LN maintains an advantage over incumbent platforms under substrate clamping, reinforcing the suitability of P3F X-cut LN for bi-directional PMUT operation in both bulk and thin-film platforms.
@@
The chosen diaphragm geometry introduces new key design parameters in addition to those reported by Lu et al., shown in Fig. 2 (b) [26]. Namely , and . is defined as the ratio of the ellipse’s minor axis (), to its major axis (), i.e., = . The parameter defines the transverse electrode offset, optimized to match the transverse energy distribution in the diaphragm while also providing space for routing to contact pads.
@@
The operating frequency was chosen to demonstrate a versatile design framework rather than targeting a specific application. The provided design parameters can be linearly scaled to achieve a resonance frequency in the lower kHz regime to leverage low attenuation, or higher MHz frequencies for superior axial resolution [51, 52].
@@ 3.3 Measurements and Post-processing
To mitigate performance degradation, the packaged device was thermally annealed in air using a Silicon Nitride (SiN) resistive heater. The temperature was ramped at approximately 25 ∘C per minute to 400 ∘C, held for 1 h, and then ramped back to room temperature. The anneal was performed in an open-air environment to avoid potential contamination from the organic epoxy used in the package.
@@
Following measurement validation and post-processing, the LN PMUT demonstrated high performance. Despite a lower resonance frequency and smaller than predicted, the device exhibited excellent efficiency. Given its mechanically robust geometry, the packaged device and the SiN heater were used to assess device performance at elevated temperatures.
@@ 4.1 Post-measurement Simulations
Post-measurement simulations were utilized to evaluate the dependence of PMUT transmit efficiency on the active layer thickness, which remains its primary limiting factor in the present geometry. Fig. 21 shows the optimized design scaled linearly with respect to the piezoelectric film thickness, illustrating varying performance at 1 MHz. These results indicate that transmit efficiency in LN PMUTs can be significantly enhanced. For instance, a 2 m bimorph active layer enhances from 0.354 nm/V to 1.14 nm/V. Given the close agreement between simulation and measurement, Fig. 21 enables a meaningful comparison with the SoA. For reference, a 3 m thick air-coupled PZT PMUT can obtain a of 4.87 nm/V at 1 MHz [56]. As a result, although high- materials, such as PZT, outperform LN in transmit efficiency at comparable compliance levels, LN provides sufficient actuation capability for high sound pressure level (SPL) operation while featuring high receive sensitivity at the same time.
@@ 4.2 Receiving Performance
For the measured 65 nm/V peak actuation displacement at a of 200, a of 160 pm/Pa is simulated. Accordingly, the resulting open-circuit sensitivity () is 2.46 mV/Pa at 775 kHz. To validate the calculated value, we compare it to the open-circuit voltage obtained directly from the FEA, shown in Fig. 22. The simulation confirms the calculated estimate, showing 2.4 mV/Pa at 775 kHz. Additionally, Fig. 22 shows that the device achieves a peak open-circuit voltage at its parallel resonance, reaching up to 38 mV/Pa. As in the case of transmit efficiency, we decouple sensitivity from by utilizing the normalized open-circuit sensitivity (). Hence, we obtain a of 12 V/Pa. Nonetheless, sensitivity is highly influenced by the interfacing circuitry, and its effect is most pronounced at the system level. However, these results cannot be experimentally validated due to the electrode failure seen in Fig. 19, motivating further studies of bimorph LN-based sensors.
While external factors play a large role in receiving performance, a comparison with incumbent PMUT platforms highlights the LN PMUT’s high voltage sensitivity. State-of-the-art 36% ScAlN-based air-coupled PMUT shows an of 1 mV/Pa at 60 kHz via a pulse-echo measurement [58]. On the other hand, its air-coupled PZT counterpart for long-range detection shows 2 mV/Pa at 40 kHz through a pitch-catch experiment [1]. The performance of these ferroelectric platforms is further enhanced by incorporating a DC bias. For example, the sensitivity in a 15% ScAlN PMUT is increased from 6.43 mV/Pa in air at 150 kHz to 9.67 mV/Pa with an 80 V bias [59]. As an alternative, PMUTs can also utilize a dual-electrode structure to improve receiving performance. For instance, [60] shows 4.4 mV/Pa at 55 kHz in 30% ScAlN. By comparison, the bimorph LFE LN PMUT achieves a comparable while operating at a higher frequency, showing potential for a highly sensitive, robust PMUT platform.
@@ 4.3 Remaining Limitations and Future Development
The remaining parasitic feedthrough observed in the post-anneal electrical measurements (Fig. 14) is critical for future LN sensor development. Although LN intrinsically exhibits low dielectric loss, parasitic feedthrough acts as a major bottleneck, limiting achievable performance. While this effect was partially mitigated by annealing, residual parasitics constrain the usable sensitivity, particularly near the parallel resonance. While thermal treatment appears beneficial, the fundamental origin of the feedthrough remains unresolved and warrants further studies.
Beyond addressing parasitic effects, the thickness scalability of the P3F LN platform presents additional opportunities for optimization. The present structure features a thick P3F piezoelectric layer to prioritize robustness and voltage sensitivity, however, Fig. 21 shows that reducing film thickness can substantially enhance transmit efficiency by increasing membrane compliance. P3F LN supports this flexibility, with high-quality P3F films demonstrated at thicknesses ranging from 100 nm to 300 m [39, 40, 61, 62]. As a result, future P3F LN PMUT implementations will focus on thickness scaling to maximize transmit efficiency while maintaining sufficient mechanical integrity for required operating conditions.
@@ Acknowledgment
The authors would like to thank Dr. Todd Bauer and Dr. Kwok-Keung Law for helpful discussions. The authors would also like to thank Dr. Bichoy Bahr, Dr. Udit Rawat, and Dr. Yao Yu from the Texas Instruments Kilby Labs for valuable feedback.